# The supermassive black hole population from seeding via collisions in Nuclear Star Clusters

### v1

@@
The coexistence of nuclear star clusters (NSCs) and supermassive black holes (SMBHs) in galaxies with stellar masses $\sim 10^{10}\leavevmode\nobreak$, the scaling relations between their properties and properties of the host galaxy (e.g., $M_{\rm NSC}^{\rm stellar}-M_{\rm galaxy}^{\rm stellar}$, $M_{\rm BH}-M_{\rm galaxy}^{\rm stellar}$), and the fact that NSCs seem to take on the role of SMBHs in less massive galaxies and vice versa in the more massive ones, suggest that the origin of NSCs and SMBHs is related. In this study, we implement an ’in-situ’ NSC formation scenario, where NSCs are formed in the center of galaxies due to star formation in the accumulated gas. We explore the impact of the free parameter $A_{\rm res}$ which regulates the amount of gas transferred to the NSC reservoir, playing a crucial role in shaping the cluster’s growth. Simultaneously, we include a BH seed formation recipe based on stellar collisions within NSCs in the Semi-Analytical Model (SAM) Galacticus to explore the resulting population of SMBHs. We determine the parameter space of the NSCs that form a BH seed and find that in initially more compact NSCs the formation of these BH seeds is more favorable, leading to the formation of light, medium and heavy BH seeds which finally reach masses up to $\sim 10^{9}$ M and is comparable with the observed SMBH mass function at masses above $10^{8}$ M. Additionally, we compare the resulting population of NSCs with a derived NSC mass function from the stellar mass function of galaxies from the GAMA survey at $z<0.06$ finding a well agreement in shape terms. We also find a considerable overlap in the observed scaling relations between the NSC mass, and the host galaxy stellar mass and velocity dispersion which is independent of the value of $A_{\rm res}$. However, the chi-square analysis suggests that the model requires further refinement to achieve better quantitative agreement.
@@ 1 Introduction
Observations show that NSCs and/or SMBHs are nearly ubiquitous in the centers of galaxies (Seth et al., 2008a; Graham & Spitler, 2009; Neumayer et al., 2011; Neumayer & Walcher, 2012; Nguyen et al., 2019). Although the coexistence of both objects is observed in almost all galaxies with stellar masses $M_{\mathrm{galaxy}}^{\mathrm{stellar}}\simeq 10^{10}\leavevmode\nobreak$ (Filippenko & Ho, 2003; González Delgado et al., 2008; Seth et al., 2008b; Neumayer et al., 2020), more massive galaxies tend to host only SMBHs (Kormendy & Ho, 2013), while less massive galaxies preferentially tend to host only NSCs. Sometimes therefore the term central massive object (CMO) is used when both types of objects are present in the host galaxy (Ferrarese et al., 2006).
It is interesting that NSCs and SMBHs follow similar scaling relations with the host-galaxy properties (see e.g., Wehner & Harris, 2006; Georgiev et al., 2016) suggesting that both undergo similar physical process in their formation and evolution. By way of illustration, previous works explored the $M_{\mathrm{NSC}}^{\mathrm{stellar}}-M_{\mathrm{galaxy}}^{\mathrm{stellar}}$ correlation (Balcells et al., 2003; Scott & Graham, 2013; den Brok et al., 2015; Sánchez-Janssen et al., 2019; Georgiev et al., 2016), which is also observed for SMBHs as $M_{\bullet}-M_{\mathrm{galaxy}}^{\mathrm{stellar}}$ (Ding et al., 2020; Tanaka et al., 2024), and other correlations between $M_{\mathrm{NSC}}^{\mathrm{stellar}}$, $M_{\bullet}$ and host galaxy luminosity or stellar velocity dispersion (Seth et al., 2008a; Gültekin et al., 2009a; Kormendy & Ho, 2013).
@@
Two principal models have been proposed to explain the formation of NSCs in the center of galaxies: (i) infall of globular clusters into the nucleus due to dynamical friction and consecutive mergers in the center, building up a NSC (Tremaine et al., 1975; Capuzzo-Dolcetta, 1993; Agarwal & Milosavljević, 2011; Antonini, 2013; Gray et al., 2024), thereby explaining the deficit of massive globular clusters (GCs) in the inner part of the nucleated early-type galaxies (Lotz et al., 2001; Capuzzo-Dolcetta & Mastrobuono-Battisti, 2009); (ii) in-situ star formation in the galactic nucleus, which happens when gas reaches the center of the galaxy, and NSCs are directly formed there (Loose et al., 1982; Antonini et al., 2015).
@@
In this work, we study the expected population of SMBHs from the collision-based BH seeding scenario utilizing the SAM for galaxy formation and evolution Galacticus (Benson, 2012). The advantage of SAMs is the fast exploration of the parameter space (Henriques et al., 2009; Benson & Bower, 2010; Bower et al., 2010) with low computational cost compared to N-Body and/or hydrodynamical simulations. Of course, the degree of approximation depends on the implemented physics, but overall a good agreement can be achieved between semi-analytic and N-body/hydrodynamical models (Benson et al., 2001; Hirschmann et al., 2012; Côté et al., 2018; Mitchell et al., 2018). In Sect. 2, we describe the implementation and the methodology of our study, where in subsection 2.1 we describe the NSC model implemented in Galacticus and in subsection 2.2 the implementation of the BH seed formation scenario. We compare our results with the observed data and summarize our findings in Sect. 3. A final discussion and conclusions are provided in Sect. 4.
@@ 2.1 Nuclear Star Cluster formation model
We assume an in-situ star formation scenario in the gas accumulated in the center of the galaxies, which is particularly well-justified for galaxies with stellar masses above $3\times 10^{11}\leavevmode\nobreak$ (Antonini et al., 2015). The gas mass accumulation rate in the reservoir is correlated to the star formation rate in the bulge (Granato et al., 2004; Haiman et al., 2004; Neumayer et al., 2011; Lapi et al., 2014; Antonini et al., 2015) and is given by
@@
The star formation rate (SFR) of the NSC is assumed to follow the law of Krumholz et al. (2009). The SFR happens in a ”quiescent” mode and takes place on a timescale $t_{\mathrm{SF}}$ as described by Krumholz et al. (2009); Sesana et al. (2014); Antonini et al. (2015), involving a fraction of cold gas ($f_{c}$) available for star formation:
@@ 2.2.2 Implementation of the black hole seeding and its evolution in Galacticus
Table 1: Initial parameters for our semi-analytic model. Column (1) corresponds to the name of the model, column (2) is the free parameter regulating the transfer of gas to the NSC gas reservoir, column (3) is the efficiency parameter describing the fraction of the stellar mass converted into a BH seed once the cluster exceeds the critical mass, column (4) is the efficiency parameter described in Sect. 2.2.1, and column (5) denotes the minimum NSC mass adopted in the respective model.
@@ 3.1 Nuclear star clusters and host galaxies
We compare the model predictions with the observed relations of NSC clusters and their host galaxies as reported by Neumayer et al. (2020). Within Table 5 in our appendix, we have extended their Table 3 including the velocity dispersion of the host galaxy from the literature. By incorporating these data points in Fig. 1, we find that our models overlap considerably with the observed population. The chi-squared value is reported in each panel. Model D1 best represents the observations. The values obtained suggest that the model is still somewhat oversimplified, as we neglect the growth via mergers with Globular Clusters, and even the in-situ channel itself may have an environmental dependence. An improvement of the model thus very likely requires an improved understanding of NSC formation itself. We further conclude that the observed population very likely does not correspond to one particular model set with a fixed value of $A_{\rm res}$, but rather the value of $A_{\rm res}$ may depend on the host galaxy and on its specific efficiency of transporting gas into the center, which potentially could be influenced by their galaxy type, the formation history as well as through effects from their environment.
@@ 3.2 Nuclear Star Cluster mass function
In general, we found that there is a shift of the NSC mas function derived from the galaxy stellar mass function (model A1 in Fig. 2), predicting less massive NSCs. This could suggest that the free parameter which regulates the gas transfer rate ($A_{\mathrm{res}}$) in the ’in-situ’ scenario is not constant over the time and may depend on environmental factors.
@@
Although no single model accurately reproduces the NSC mass function derived from the galaxy stellar mass function across all mass ranges, we highlight that models D1, E1, and F1 provide a good fit to the high-mass end of the NSC mass function. This aligns with the current understanding of NSC formation mechanisms in massive galaxies, where the in-situ formation scenario is typically invoked.
@@ 3.3 Nuclear Star Clusters at critical mass
Table 2: Range of BH masses formed in our semi-analytical models. (1) Name of the model; (2) Minimum and maximum mass of the BH seed formed in solar masses.
@@
Figure 6: Same as Fig. 5 but for models D7 (brown dots) and D4 (teal dots), with $M_{\mathrm{threshold}}=10^{4}$ M and $\epsilon_{r}=0.5$ and $0.1$, respectively, as listed in Table  1. Models D5 and D6 do not form any black hole seed as the conditions for seeding ($M_{\mathrm{crit}}\leq M_{\mathrm{NSC}}^{\mathrm{stellar}}$ and $M_{\mathrm{threshold}}\geq M_{\mathrm{NSC}}^{\mathrm{stellar}}$) are not fulfilled for these models.
@@ 3.4 Black hole mass function
For a detailed quantitative analysis of the over- or underestimation of the Black Hole Mass Function (BHMF) in our models, we kindly refer readers to our Appendix B.
@@ 3.5 Nuclear star clusters, supermassive black holes and host galaxies
$\log{\left(\frac{M_{\mathrm{BH}}}{M_{\mathrm{NSC}}^{\mathrm{stellar}}}\right)}=c_{1}+c_{2}\log{\left(M_{\mathrm{galaxy}}^{\mathrm{stellar}}\right)},$ (23)
@@
We find that there is a scaling relation between the stellar mass of the NSC and the stellar mass of the galaxy in all our models as well as in the observed population. In our models, galaxies with stellar masses between $10^{9}$ M and $10^{11}$ M show a scatter in the stellar mass of the NSC in galaxies hosting a SMBH. A possible explanation for this is the fact of the SMBH accreting material from the NSC gas reservoir, removing a fraction of the available gas for star formation in the NSC. This loss of gas directly affects the star formation rate of the NSC as $\dot{M}_{\mathrm{NSC}}^{\mathrm{stellar}}\propto M_{\mathrm{NSC}}^{\mathrm{gas}}$, suggesting that BH suppress the in-situ star formation in NSCs.
@@
Table 3: Values of the best fit parameters $c_{1}$ and $c_{2}$. (1) Name of the model. (2, 3) Intercept coefficient and slope coefficient appearing in Eq. 23 with the respective ±$\pm 1\sigma$ value.
@@ 4 Discussion and conclusions
In this work we employed the semi-analytic model Galacticus developed by Benson (2012) to explore the seeding of SMBHs via collisions in NSCs, considering the framework proposed by Escala (2021), where SMBHs are formed from (partially) failed NSCs with average collision timescale shorter than the age of the system, thus implying the efficient formation of a central massive object via runaway collisions. Particularly, we employ here the concept of the critical mass scale, for which the average timescale for collisions within the NSC (of a given radius) is equal to the age of the cluster, that was introduced by Vergara et al. (2023) in order to numerically test this framework (with successful results) and later extended its validity to a range of stellar systems (globular clusters and ultra-compact dwarf galaxies, in addition to NSCs; Vergara et al., 2024). The formation of NSCs themselves is modeled here via the in situ mechanism, which is adequate for the more massive galaxy population and thus for the formation of SMBH in more massive galaxies (Antonini & Perets, 2012; Antonini et al., 2015).
@@
The overestimation of the population may indicate that the ’in-situ’ star formation scenario is enough to explain the population of the most massive NSCs but it is not the main mechanism in the low mass regime, suggesting that the invocation of other formation mechanisms such as the migration of globular clusters (e.g., Tremaine et al., 1975; Capuzzo-Dolcetta, 1993) is also required and the value of $A_{\mathrm{res}}$ is not constant for all galaxies, depending on environmental factors. Some previous studies suggest a correlation between the morphological type of the host galaxy and the formation scenario of the NSC, to be specific, globular cluster infall scenario is related to early type galaxies (Tremaine et al., 1975; Capuzzo-Dolcetta, 1993), while in massive late type galaxies the in-situ formation dominates the growth of the NSC (Pinna et al., 2021; Hannah et al., 2021). In our models, we assume that the gas transfer to the nuclear reservoir is related to the star formation in the spheroidal component of the galaxies. This is consistent with galaxies which typically have abundant gas and provides a rich gas reservoir for in-situ star formation within the nucleus of the galaxies. As our model does not account for the population of NSCs formed in bulgeless galaxies or those without a significant spheroidal component, where dynamical processes are the main formation mechanism, we are missing part of the population of NSCs formed via the cluster infall scenario. This scenario is particularly predominant in early-type galaxies, where NSCs are more likely to form through the accumulation and merging of pre-existing globular clusters that migrate to the center over time as previously explored by Capuzzo-Dolcetta (1993) in a triaxial galaxy model, N-body simulations (Milosavljević, 2004) and more recently in self-consistent numerical simulations (Bekki, 2010).
@@
In general, all our models are able to produce BH seeds, except those with $\epsilon_{r}=1$, and only few of these form in models with $\epsilon_{r}=0.5$. This may suggest that also in the observed population, the mechanism operates preferably in the more compact part of the population, or possibly also at early stages where the NSCs were still more compact. Overall, the models are able to produce a range of BHs from light seeds ($\sim 10^{2}$ M) to heavy seeds ($\sim 10^{5}$ M). The formation of the most massive seeds ($\sim 10^{5}$ M) is independent of the mass threshold we employ for the NSCs, but only happens in models where $\epsilon_{r}=0.1$. Furthermore, those seeds are formed in systems with ages of $\gtrsim 4$ Gyrs and velocity dispersions higher than $70$ km s-1. The formation of these massive seeds however happens relatively late, that is at redshifts $z\sim 0.35$ within the halos modeled here. We cannot exclude that in some cases, this may also happen early, and in this respect it might be important to study the halo merger history of the first quasars forming around $z\sim 6$, to obtain more insight into their specific histories and to understand whether this channel might be contributing there or not. In this context, also the presence of gas within the NSCs might be even more relevant, and it could be important to include the extension of the critical mass concept in the presence of gas as recently proposed by Vergara et al. (2024). In addition, it is clear that other seeding scenarios may also contribute to the general population, and the interaction between different seeding mechanisms (i.e., due to mergers) needs to be explored. This includes for example seeding via the direct collapse (Bromm & Loeb, 2003; Koushiappas et al., 2004; Wise et al., 2008; Latif et al., 2013, 2022) as well as mixed scenarios considering the interaction of collisions and accretion (Boekholt et al., 2018; Chon & Omukai, 2020; Schleicher et al., 2022; Reinoso et al., 2023; Schleicher et al., 2023).
@@
The models reproduce the observed mass functions of NSCs and SMBHs, with a good match in the main mass range relevant for supermassive black hole formation. The predicted scaling relations also show considerable overlap with observed data. Around BH masses of $10^{7}$ M, our model overpredicts the observed population (constructed from the sample of Vika et al. 2009) by roughly a factor of $10$, while at higher masses we tend to underpredict the observed population. However, this of course may change when additional black hole formation mechanisms are being considered, and some of the BH seeds predicted by our model may also merge with seeds produced via different channels (Sassano et al., 2021; Trinca et al., 2022). Nonetheless, the comparison shows that the collision-based channel may contribute in a relevant way to the overall population. This could include a mechanism that is related but still different, for example, the formation of massive black holes in dark cores that are formed in the center of NSCs (e.g., Davies et al., 2011; Lupi et al., 2014; Kroupa et al., 2020; Gaete et al., 2024). We expect that more light can be shed on the seeds of SMBHs with upcoming observations both with the JWST and the Extremely Large Telescope (ELT) [^4] Webpage ELT: https://elt.eso.org/, and together with the emission of gravitational waves detected by current interferometers, such as the LIGO [^5] Webpage LIGO: https://www.ligo.caltech.edu /Virgo [^6] Webpage Virgo: https://www.virgo-gw.eu/KAGRA [^7] Webpage KRAGA:  https://gwcenter.icrr.u-tokyo.ac.jp/ collaboration (Abbott et al., 2024), and in the future with LISA (Amaro-Seoane et al., 2017) and Einstein Telescope [^8] Webpage ET: https://www.et-gw.eu (Punturo et al., 2010) both methods in the local and the high-redshift Universe.

### v2

@@
The coexistence of nuclear star clusters (NSCs) and supermassive black holes (SMBHs) in galaxies with stellar masses $\sim 10^{10}\leavevmode\nobreak$, the scaling relations between their properties and properties of the host galaxy (e.g., $M_{\rm NSC}^{\rm stellar}-M_{\rm galaxy}^{\rm stellar}$, $M_{\rm BH}-M_{\rm galaxy}^{\rm stellar}$), and the fact that NSCs seem to take on the role of SMBHs in less massive galaxies and vice versa in the more massive ones, suggest that the origin of NSCs and SMBHs is related. In this study, we implement an ‘in-situ’ NSC formation scenario, where NSCs are formed in the center of galaxies due to star formation in the accumulated gas. We explore the impact of the free parameter $A_{\rm res}$ which regulates the amount of gas transferred to the NSC reservoir, playing a crucial role in shaping the cluster’s growth. Simultaneously, we include a BH seed formation recipe based on stellar collisions within NSCs in the Semi-Analytical Model (SAM) Galacticus to explore the resulting population of SMBHs. We determine the parameter space of the NSCs that form a BH seed and find that in initially more compact NSCs the formation of these BH seeds is more favorable, leading to the formation of light, medium and heavy BH seeds which finally reach masses up to $\sim 10^{9}$ M and is comparable with the observed SMBH mass function at masses above $10^{8}$ M. Additionally, we compare the resulting population of NSCs with a derived NSC mass function from the stellar mass function of galaxies from the GAMA survey at $z<0.06$ finding a well agreement in shape terms. We also find a considerable overlap in the observed scaling relations between the NSC mass, the stellar mass of the host galaxy, and the velocity dispersion, which is independent of the value of $A_{\rm res}$. However, the chi-square analysis suggests that the model requires further refinement to achieve better quantitative agreement.
@@ 1 Introduction
Observations show that nuclear star clusters (NSCs) and/or supermassive black holes (SMBHs) are nearly ubiquitous in the centers of galaxies (Seth et al., 2008a; Graham & Spitler, 2009; Neumayer et al., 2011; Neumayer & Walcher, 2012; Nguyen et al., 2019). Although the coexistence of both objects is observed in almost all galaxies with stellar masses $M_{\mathrm{galaxy}}^{\mathrm{stellar}}\simeq 10^{10}\leavevmode\nobreak$ (Filippenko & Ho, 2003; González Delgado et al., 2008; Seth et al., 2008b; Neumayer et al., 2020), more massive galaxies tend to host only SMBHs (Kormendy & Ho, 2013), while less massive galaxies preferentially tend to host only NSCs. Sometimes therefore the term central massive object (CMO) is used when both types of objects are present in the host galaxy (Ferrarese et al., 2006).
It is interesting that NSCs and SMBHs follow similar scaling relations with the host-galaxy properties (see e.g., Wehner & Harris, 2006; Georgiev et al., 2016) suggesting that both undergo similar physical process in their formation and evolution. By way of illustration, previous works explored the $M_{\mathrm{NSC}}^{\mathrm{stellar}}-M_{\mathrm{galaxy}}^{\mathrm{stellar}}$ correlation (Balcells et al., 2003; Scott & Graham, 2013; den Brok et al., 2015; Sánchez-Janssen et al., 2019; Georgiev et al., 2016), which is also observed for SMBHs as $M_{\rm BH}-M_{\mathrm{galaxy}}^{\mathrm{stellar}}$ (Ding et al., 2020; Tanaka et al., 2024), and other correlations between $M_{\mathrm{NSC}}^{\mathrm{stellar}}$, $M_{\rm BH}$ and host galaxy luminosity or stellar velocity dispersion (Seth et al., 2008a; Gültekin et al., 2009a; Kormendy & Ho, 2013).
@@
Two principal models have been proposed to explain the formation of NSCs in the center of galaxies: (i) infall of globular clusters into the nucleus due to dynamical friction and consecutive mergers in the center, building up a NSC (Tremaine et al., 1975; Capuzzo-Dolcetta, 1993; Agarwal & Milosavljević, 2011; Antonini, 2013; Gray et al., 2024), thereby explaining the deficit of massive globular clusters (GCs) in the inner part of the nucleated early-type galaxies (Lotz et al., 2001; Capuzzo-Dolcetta & Mastrobuono-Battisti, 2009); (ii) ‘in-situ’ star formation in the galactic nucleus, which happens when gas reaches the center of the galaxy, and NSCs are directly formed there (Loose et al., 1982; Antonini et al., 2015).
@@
In this work, we study the expected population of SMBHs from the collision-based BH seeding scenario utilizing the semi-analytic model (SAM) for galaxy formation and evolution Galacticus (Benson, 2012). The advantage of SAMs is the fast exploration of the parameter space (Henriques et al., 2009; Benson & Bower, 2010; Bower et al., 2010) with low computational cost compared to N-Body and/or hydrodynamical simulations. Of course, the degree of approximation depends on the implemented physics, but overall a good agreement can be achieved between semi-analytic and N-body/hydrodynamical models (Benson et al., 2001; Hirschmann et al., 2012; Côté et al., 2018; Mitchell et al., 2018). In Sect. 2, we describe the implementation and the methodology of our study, where in subsection 2.1 we describe the NSC model implemented in Galacticus and in subsection 2.2 the implementation of the BH seed formation scenario. We compare our results with the observed data and summarize our findings in Sect. 3. A final discussion and conclusions are provided in Sect. 4.
@@ 2.1 Nuclear Star Cluster formation model
We assume an ‘in-situ’ star formation scenario in the gas accumulated in the center of the galaxies, which is particularly well-justified for galaxies with stellar masses above $3\times 10^{11}\leavevmode\nobreak$ (Antonini et al., 2015). The gas mass accumulation rate in the reservoir is correlated to the star formation rate in the bulge (Granato et al., 2004; Haiman et al., 2004; Neumayer et al., 2011; Lapi et al., 2014; Antonini et al., 2015) and is given by
@@
The star formation rate (SFR) of the NSC is assumed to follow the law of Krumholz et al. (2009). The SFR happens in a ‘quiescent’ mode and takes place on a timescale $t_{\mathrm{SF}}$ as described by Krumholz et al. (2009); Sesana et al. (2014); Antonini et al. (2015), involving a fraction of cold gas ($f_{c}$) available for star formation:
@@ 2.2.2 Implementation of the black hole seeding and its evolution in Galacticus
Table 1: Initial parameters for our semi-analytic model.
@@
444(1) Corresponds to the name of the model;(2) free parameter regulating the transfer of gas to the NSC gas reservoir; (3) efficiency parameter describing the fraction of the stellar mass converted into a BH seed once the cluster exceeds the critical mass; (4) efficiency parameter described in Sect. 2.2.1; (5) denotes the minimum NSC mass adopted in the respective model.
@@ 3.1 Nuclear star clusters and host galaxies
We compare the model predictions with the observed relations of NSC clusters and their host galaxies as reported by Neumayer et al. (2020). Within Table 5 in our appendix, we have extended their Table 3 including the velocity dispersion of the host galaxy from the literature. By incorporating these data points in Fig. 1, we find that our models overlap considerably with the observed population. The chi-squared value is reported in each panel. Model D1 best represents the observations. The values obtained suggest that the model is still somewhat oversimplified, as we neglect the growth via mergers with Globular Clusters, and even the ‘in-situ’ channel itself may have an environmental dependence. An improvement of the model thus very likely requires an improved understanding of NSC formation itself. We further conclude that the observed population very likely does not correspond to one particular model set with a fixed value of $A_{\rm res}$, but rather the value of $A_{\rm res}$ may depend on the host galaxy and on its specific efficiency of transporting gas into the center, which potentially could be influenced by their galaxy type, the formation history as well as through effects from their environment.
@@ 3.2 Nuclear Star Cluster mass function
In general, we found that there is a shift of the NSC mas function derived from the galaxy stellar mass function (model A1 in Fig. 2), predicting less massive NSCs. This could suggest that the free parameter which regulates the gas transfer rate ($A_{\mathrm{res}}$) in the ‘in-situ’ scenario is not constant over the time and may depend on environmental factors.
@@
Although no single model accurately reproduces the NSC mass function derived from the galaxy stellar mass function across all mass ranges, we highlight that models D1, E1, and F1 provide a good fit to the high-mass end of the NSC mass function. This aligns with the current understanding of NSC formation mechanisms in massive galaxies, where the ‘in-situ’ formation scenario is typically invoked.
@@ 3.3 Nuclear Star Clusters at critical mass
Table 2: Range of BH masses formed in our semi-analytical models.
@@
555(1) Name of the model; (2) minimum and maximum mass of the BH seed formed in solar masses.
@@
Figure 6: Same as Fig. 5 but for models D7 (brown dots) and D4 (teal dots), with $M_{\mathrm{threshold}}=10^{4}$ M and $\epsilon_{r}=0.5$ and $0.1$, respectively, as listed in Table  4. Models D5 and D6 do not form any black hole seed as the conditions for seeding ($M_{\mathrm{crit}}\leq M_{\mathrm{NSC}}^{\mathrm{stellar}}$ and $M_{\mathrm{threshold}}\geq M_{\mathrm{NSC}}^{\mathrm{stellar}}$) are not fulfilled for these models.
@@ 3.4 Black hole mass function
For a detailed quantitative analysis of the over- or underestimation of the BHMF in our models, we kindly refer readers to our Appendix B.
@@ 3.5 Nuclear star clusters, supermassive black holes and host galaxies
$\log{\left(\frac{M_{\mathrm{BH}}}{M_{\mathrm{NSC}}^{\mathrm{stellar}}}\right)}=c_{1}+c_{2}\log{\left(\frac{M_{\mathrm{galaxy}}^{\mathrm{stellar}}}{1\leavevmode\nobreak\ {\rm M}_{\odot}}\right)},$ (23)
@@
We find that there is a scaling relation between the stellar mass of the NSC and the stellar mass of the galaxy in all our models as well as in the observed population. In our models, galaxies with stellar masses between $10^{9}$ M and $10^{11}$ M show a scatter in the stellar mass of the NSC in galaxies hosting a SMBH. A possible explanation for this is the fact of the SMBH accreting material from the NSC gas reservoir, removing a fraction of the available gas for star formation in the NSC. This loss of gas directly affects the star formation rate of the NSC as $\dot{M}_{\mathrm{NSC}}^{\mathrm{stellar}}\propto M_{\mathrm{NSC}}^{\mathrm{gas}}$, suggesting that BH suppress the ‘in-situ’ star formation in NSCs.
Table 3: Values of the best fit parameters $c_{1}$ and $c_{2}$.
@@
666 (1) Name of the model; (2, 3) intercept coefficient and slope coefficient appearing in Eq. 23 with the respective ±$\pm 1\sigma$ value.
@@ 4 Discussion and conclusions
In this work we employed the semi-analytic model Galacticus developed by Benson (2012) to explore the seeding of SMBHs via collisions in NSCs, considering the framework proposed by Escala (2021), where SMBHs are formed from (partially) failed NSCs with average collision timescale shorter than the age of the system, thus implying the efficient formation of a central massive object via runaway collisions. Particularly, we employ here the concept of the critical mass scale, for which the average timescale for collisions within the NSC (of a given radius) is equal to the age of the cluster, that was introduced by Vergara et al. (2023) in order to numerically test this framework (with successful results) and later extended its validity to a range of stellar systems (globular clusters and ultra-compact dwarf galaxies, in addition to NSCs; Vergara et al., 2024). The formation of NSCs themselves is modeled here via the ‘in-situ’ mechanism, which is adequate for the more massive galaxy population and thus for the formation of SMBH in more massive galaxies (Antonini & Perets, 2012; Antonini et al., 2015).
@@
The overestimation of the population may indicate that the ‘in-situ’ star formation scenario is enough to explain the population of the most massive NSCs but it is not the main mechanism in the low mass regime, suggesting that the invocation of other formation mechanisms such as the migration of globular clusters (e.g., Tremaine et al., 1975; Capuzzo-Dolcetta, 1993) is also required and the value of $A_{\mathrm{res}}$ is not constant for all galaxies, depending on environmental factors. Some previous studies suggest a correlation between the morphological type of the host galaxy and the formation scenario of the NSC, to be specific, globular cluster infall scenario is related to early type galaxies (Tremaine et al., 1975; Capuzzo-Dolcetta, 1993), while in massive late type galaxies the ‘in-situ’ formation dominates the growth of the NSC (Pinna et al., 2021; Hannah et al., 2021). In our models, we assume that the gas transfer to the nuclear reservoir is related to the star formation in the spheroidal component of the galaxies. This is consistent with galaxies which typically have abundant gas and provides a rich gas reservoir for ‘in-situ’ star formation within the nucleus of the galaxies. As our model does not account for the population of NSCs formed in bulgeless galaxies or those without a significant spheroidal component, where dynamical processes are the main formation mechanism, we are missing part of the population of NSCs formed via the cluster infall scenario. This scenario is particularly predominant in early-type galaxies, where NSCs are more likely to form through the accumulation and merging of pre-existing globular clusters that migrate to the center over time as previously explored by Capuzzo-Dolcetta (1993) in a triaxial galaxy model, N-body simulations (Milosavljević, 2004) and more recently in self-consistent numerical simulations (Bekki, 2010).
@@
In general, all our models are able to produce BH seeds, except those with $\epsilon_{r}=1$, and only few of these form in models with $\epsilon_{r}=0.5$. This may suggest that also in the observed population, the mechanism operates preferably in the more compact part of the population, or possibly also at early stages where the NSCs were still more compact. Overall, the models are able to produce a range of BHs from light seeds ($\sim 10^{2}$ M) to heavy seeds ($\sim 10^{5}$ M). The formation of the most massive seeds ($\sim 10^{5}$ M) is independent of the mass threshold we employ for the NSCs, but only happens in models where $\epsilon_{r}=0.1$. Furthermore, those seeds are formed in systems with ages of $\gtrsim 4$ Gyrs and velocity dispersions higher than $70$ km s-1. The formation of these massive seeds however happens relatively late, that is at redshifts $z\sim 0.35$ within the halos modeled here. We cannot exclude that in some cases, this may also happen early, and in this respect it might be important to study the halo merger history of the first quasars forming around $z\sim 6$, to obtain more insight into their specific histories and to understand whether this channel might be contributing there or not. In this context, also the presence of gas within the NSCs might be even more relevant, and it could be important to include the extension of the critical mass concept in the presence of gas as recently proposed by Vergara et al. (2024). In addition, it is clear that other seeding scenarios may also contribute to the general population, and the interaction between different seeding mechanisms (i.e., due to mergers) needs to be explored. This includes, for example, seeding via the direct collapse (Bromm & Loeb, 2003; Koushiappas et al., 2004; Wise et al., 2008; Latif et al., 2013, 2022) as well as mixed scenarios considering the interaction of collisions and accretion (Boekholt et al., 2018; Chon & Omukai, 2020; Schleicher et al., 2022; Reinoso et al., 2023; Schleicher et al., 2023).
The models successfully reproduce the shape of the observed mass functions of NSCs and SMBHs, showing a strong match within the mass range relevant to SMBH formation. However, they overpredict the NSC and BH mass functions at lower masses, likely due to an excess (by a factor of $\sim 10$) in the number of galaxies predicted by Galacticus compared to the Baldry et al. (2012) sample.
Despite this discrepancy, our results demonstrate that the proposed scenario effectively forms black hole seeds capable of growing to masses up to $10^{9}$ M. A logical next step is to enhance the statistical robustness of our predictions. Unfortunately, rescaling the stellar mass function is not straightforward. One potential approach is to employ the Markov Chain Monte Carlo (MCMC) functionality built into Galacticus. While MCMC can improve statistical predictions, it requires a significant computational investment to achieve reliable results.
The predicted scaling relations also show considerable overlap with observed data. Around BH masses of $10^{7}$ M, our model overpredicts the observed population (constructed from the sample of Vika et al. 2009) by roughly a factor of $10$, while at higher masses we tend to underpredict the observed population. However, this of course may change when additional black hole formation mechanisms are being considered, and some of the BH seeds predicted by our model may also merge with seeds produced via different channels (Sassano et al., 2021; Trinca et al., 2022).
@@
Nonetheless, the comparison shows that the collision-based channel may contribute in a relevant way to the overall population. This could include a mechanism that is related but still different, for example, the formation of massive black holes in dark cores that are formed in the center of NSCs (e.g., Davies et al., 2011; Lupi et al., 2014; Kroupa et al., 2020; Gaete et al., 2024). We expect that more light can be shed on the seeds of SMBHs with upcoming observations both with the JWST and the Extremely Large Telescope (ELT) [^7] Webpage ELT: https://elt.eso.org/, and together with the emission of gravitational waves detected by current interferometers, such as the Laser Interferometer Gravitational-Wave Observatory (LIGO) [^8] Webpage LIGO: https://www.ligo.caltech.edu /Virgo [^9] Webpage Virgo: https://www.virgo-gw.eu/ Kamioka Gravitational Wave Detector (KAGRA) [^10] Webpage KRAGA:  https://gwcenter.icrr.u-tokyo.ac.jp/ collaboration (Abbott et al., 2024), and in the future with LISA (Amaro-Seoane et al., 2017) and Einstein Telescope [^11] Webpage ET: https://www.et-gw.eu (Punturo et al., 2010) both methods in the local and the high-redshift Universe.